It costs you about $3 to produce a pizza.

Last week you sold 500 pizzas for $12 each.

This week you raised your price and sold 300 pizzas for $16 each.

What price should you be charging? (round to the nearest penny)
1
Expert's answer
2017-12-11T11:46:07-0500
500=a*12+b
300=a*16+b
200=-4a
a=-50
b=1100
Q=-50P+1100
P=0.02Q+22
TR=P*Q=-0.02Q*Q+22Q
MR=-0.04Q+22
3=-0.04Q+22
Q=475
P=12.5